# Bunch-by-Bunch Prediction of Beam Transverse Position, Phase, and Length in a Storage Ring Using Neural Networks

### v1

@@ II.2 Calculating Theoretical 3D Position
$x$ $=k_{x}\cdot\frac{{amp}_{A}-{amp}_{B}-{amp}_{C}+{amp}_{D}}{{amp}_{A}+{amp}_{B}+{amp}_{C}+{amp}_{D}}$ (2)
$y$ $=k_{y}\cdot\frac{{amp}_{A}+{amp}_{B}-{amp}_{C}-{amp}_{D}}{{amp}_{A}+{amp}_{B}+{amp}_{C}+{amp}_{D}}$ (3)
@@ III.1.4 Dynamic Feature Fusion Module
$X_{Tshift}=\text{ReLU}(W_{e}\cdot Tshift+b_{e})$ (6)
$z=\alpha\cdot\text{attended}_{out}+(1-\alpha)\cdot X_{Tshift}$ (7)
@@ V.3 Longitudinal Phase Prediction Results
On the HLS-II dataset, the model achieved a performance of $R^{2}=0.973$ and MAE = 3.59 fs, with a total inference time of 1.19 seconds. The prediction results for a subset of samples are visualized in Fig. 10. In addition to demonstrating a close alignment between the predicted and true values, the two subfigures collectively reveal that the predominant errors are concentrated near the zero-crossings and the peaks/troughs of the phase trajectory. This is likely because the relationship between the BPM waveform and the beam phase is highly non-linear at these turning points, making precise predictions more challenging. Nevertheless, given the total phase range of -120 fs to 120 fs, the achieved MAE still represents a relatively small error.
@@
The model also exhibited excellent performance on the SSRF dataset, achieving an $R^{2}$ of 0.9886 and an MAE of 0.5040 fs over 387 samples. The inference was completed in 0.10 seconds. Fig. 11 illustrates the prediction results for SSRF, showing a similarly high fidelity in tracking the phase evolution. Notably, for SSRF, the error distribution also followed a similar pattern, with slight inaccuracies more prevalent at the phase extrema and zero-crossings. The consistent performance across two distinct facilities underscores the generalizability and robustness of the proposed phase prediction architecture. This high-fidelity phase prediction enables direct application in advanced beam dynamics studies, such as the experimental analysis of beam loading effects in storage rings [26].

### v2

@@ II.2 Calculating Theoretical 3D Position
{align}
x = k_x ⋅ampA- ampB- ampC+ ampDampA+ ampB+ ampC+ ampD
y = k_y ⋅ampA+ ampB- ampC- ampDampA+ ampB+ ampC+ ampD
@@ III.1.4 Dynamic Feature Fusion Module
{gather}
X_Tshift = ReLU(W_e ⋅Tshift + b_e)
z = αattended_out + (1 - α) ⋅X_Tshift
@@ V.3 Longitudinal Phase Prediction Results
On the HLS-II dataset, the model demonstrated high predictive accuracy, achieving an $R^{2}=0.973$ and a Mean Absolute Error (MAE) of 3.59 fs, with a total inference time of 1.19 seconds. The prediction results for a representative subset of samples are visualized in Fig. 10. As illustrated in subfigure (a), the ”True vs. Predicted Values Comparison” shows a blue solid line representing the ground truth and a red dashed line representing the model’s predictions. The two curves exhibit close alignment across the 500-sample index, indicating the model’s capacity to track the rapid oscillations in phase. This performance is further validated by the ”True vs. Predicted Values Scatter Plot” in subfigure (b), where the data points are tightly clustered along the identity line. The residuals are uniformly distributed across the operational range of -150 fs to +150 fs, demonstrating consistent predictive reliability and a lack of significant bias across the entire phase interval.
@@
The model also demonstrated exceptional performance on the SSRF dataset, achieving an $R^{2}$ of 0.9886 and an MAE of 0.5040 fs across 387 samples, with a total inference time of 0.10 seconds. The prediction results for SSRF are shown in Fig. 11, illustrating high-fidelity tracking of the phase evolution. Although a marginal increase in prediction residuals is observed near the phase extrema and zero-crossings, the model maintains high accuracy across the full signal profile. This slight deviation at turning points is likely attributable to the heightened nonlinearity in the mapping relationship between the beam position monitor (BPM) waveform and the beam phase in these specific intervals.
The model’s consistent high performance across two distinct experimental facilities confirms the strong generalizability and robustness of the proposed phase prediction architecture. Such high-fidelity phase prediction enables its direct application in advanced beam dynamics studies, such as the experimental analysis of beam loading effects in storage rings [26].